OpenStudy (anonymous):

How do I figure out step by step if this function is even, odd, or neither? f(x)=(x)/(x^2+1)

OpenStudy (amistre64):

plug in x= -x; and if you get back f(x), its even if you get back -f(x) its odd if neither, then its neither

OpenStudy (amistre64):

\[f(-x)=\frac{(-x)}{(-x)^2+1}=-\frac{x}{x^2+1}=-f(x)\]

OpenStudy (anonymous):

so.. f(-x)=(-x)/(-x^2+1)=(-x)/(-x^2+1) ?

OpenStudy (amistre64):

so far yes, but dont forget to wrap it ias (-x)^2 + 1

OpenStudy (anonymous):

so after that what do I do?

OpenStudy (amistre64):

when i first learned these, the replacing x by -x seemed wierd; so if you want to replace it by a different negative letter thats fine

OpenStudy (amistre64):

look about 4 posts up ... i worked it out

OpenStudy (anonymous):

ohh and if it comes out positive its even if it comes out negative its an odd?

OpenStudy (amistre64):

if it comes out exactly the same; its even; for example: I know x^2 is even f(-x) = (-x)^2 = x^2 ; since x^2 = x^2; f(-x) = f(x)

OpenStudy (amistre64):

i know f(x) = x is an odd function; f(-x) = -x = -f(x)

OpenStudy (anonymous):

and in my problem; its an odd function because x resulted in (-x) even though x^2+1 came back the same.. (-x)^2+1=x^2+1.. so if my problem was x^2/x^2+1... that would be an even function?

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

you want to look at the overall effect it has on the function; not just parts of it

OpenStudy (amistre64):

if the overall function: f(-x) simplifies out to f(x) its even if it simplifies out to the opposite of f(x): also known as -f(x), then its off if you cant get it into either of these, then its neither of them ...

OpenStudy (anonymous):

okay.. so for f(x)=(x^2)/(x^4+1)...f(-x)=(-x)^2/(-x)^4+1=x^2/x^4+1=even function?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

yay! thanks.

OpenStudy (amistre64):

for rational expressions, after a few of them you can develop a sense of their properties; they mimic the +- products. even/even = even odd/even = odd even/odd = odd odd/odd = even

OpenStudy (anonymous):

Okay. I'll remember that :)

OpenStudy (anonymous):

@amistre64 ... would f(x)=x|x| be and odd function.. f(-x)=(-x)|-x| =f(-x)=-x|x|

OpenStudy (amistre64):

|x| is even x is odd even * odd = odd

OpenStudy (amistre64):

and yes, the more proofy way: f(-x) = (-x)|-x| = - [x|x|] = -f(x)

OpenStudy (anonymous):

okay thank you!

OpenStudy (amistre64):

youre welcome